Reports 1-1 of 1 Clear search Modify search
Virgo Runs (O3)
chiummo, Gherardini - 10:51 Sunday 01 December 2019 (47826) Print this report
Comment to Operator Report - Night shift (47824)

Due to the HVAC failure, a large temperature excursion occurred in the laser lab. This, among other effects, has triggered a drift of the beam pointing compensated by the BPC loop with large corrections, especially on the horinzontal degrees of freedom (see first plot).

Unfortunately, after the fixing of the HVAC, the temperatures took a long time to stabilize, and are still experiencing a long transient. Nevertheless, the drift of the beam pointing had become too large to be dealt with by the correction dynamics, so I come on site to decrease the needed corrections of the BPC by acting on the usual M6-M8 picomotors. The operation has been successful (see second plot), but fine tuning will be needed once that the steady state of the temperature will be achieved at the right values.

The interferometer was able to relock up to LN3 and was set in science mode after the operation.

Images attached to this comment
Search Help
×

Warning

×